Update 3:47a: We will be engaging our vendor on why our DHCP server failed to perform as expected. If you are having connectivity issues please reboot your ONT and router. If problems persist please contact support.
Update 2:52a: We are still working to get the DHCP server for the Berkeley area up to full capacity. If you are experiencing an outage please reboot your ONT and router.
Update 1:39a: Maintenance is complete in most areas. We are currently investigating a DHCP issue in the Berkeley area which is causing service restoral for that area to take much longer than expected.
Wednesday 1/10 starting at 11:59pm, we will be performing vendor upgrades for equipment serving subsets of our Fiber customers in Northern California. Expected service interruption will be 5-10 minutes while equipment reloads. The maintenance window is until 4am. An update to this MOTD will be provided when the work is completed.